36.) If a manufacturer charges q dollars each for footballs, then he can sell 3000 - 150q footballs per week. Find a polynomial R(q) that represents the revenue for one week. Find the weekly revenue if the price is $8 for each football.
This is the answer that it is showing r(q)= -150q^2 + 3000q, 14,400
Can you show me how they got this answer please.

Answer by ankor@dixie-net.com(22740)

This football problem has been kicking around for few days.
:
If a manufacturer charges q dollars each for footballs, then he can sell 3000 - 150q footballs per week. Find a polynomial R(q) that represents the revenue for one week.
:
Revenue = price times number sold (q=price), (number sold = 3000-150q)
R(q) = q(3000-150q)
;
Multiply what's inside the brackets
R(q) = 3000q - 150q^2
or
R(q) = -150q^2 + 3000q
:
:
Find the weekly revenue if the price is $8 for each football.
:
Take the above equation, substitute 8 for q and find R(q)
R(q) = -150(8^2) + 3000(8)
:
R(q) = -150(64) + 24000
:
R(q) = -9600 + 24000
:
R(q) = $14,400
:
:
Did this explain it to you?